OpenStudy (anonymous):

[6.02] For the following system, if you isolated x in the first equation to use the Substitution Method, what expression would you substitute into the second equation? −x + 2y = 9 4x + 5y = 9 a. −9 + 2y b. −9 − 2y c. 9 − 2y d. 9 + 2y

OpenStudy (anonymous):

Okay. It's basically asking to get x by itself in the first equation. -x+2y=9 -x=9-2y x=-9+2y A.

OpenStudy (anonymous):

Wow I appreciate it, I had trouble with the problem for a long time! is there a way to vouche for you?

OpenStudy (anonymous):

No problem. You can ... thanks! :)

OpenStudy (anonymous):

How can I? would gladly do so :[)

OpenStudy (anonymous):

You just did. Good luck! (Don't forget to close the question).